For the following relation, write an equation that describes the connection between x (the first component in an ordered pair) and y (the second component in an ordered pair).




a. {(1, 1), (1, -1), (4, 2), (4, -2), (9, 3), (9, -3)}








b. {(1, 4), (2, 3), (3, 2), (4, 1), (-2, -3), (6, -1), (7, -2)}




1
Expert's answer
2022-03-01T04:40:48-0500

a.


"|y|=\sqrt{x}"

b.


"y=\begin{cases}
 -x+5&x\not=-2 \\
 -3 &x=-2
\end{cases}"